# Multimodal Representation Learning and Fusion thanks:

### v1

@@ Abstract
Index Terms:
multimodal learning, representation learning, neural architecture search, data heterogeneity, resource-efficient learning, benchmark evaluation, fusion strategies, robust multimodal models, machine learning, deep learning
@@ I Introduction
Multi-modal learning, a fast-growing research field that aims to enable machines to interpret and understand the world by combining different types of data, such as images, text, and audio [Parcalabescu et al., 2021] [Baltrušaitis et al., 2018]. It tries to use the strengths of each modality to build a more complete and detailed understanding of complex things [Akkus et al., 2023]. This kind of learning is naturally cross-disciplinary, combining knowledge from computer vision, natural language processing, speech analysis, and other fields to make machines behave more like humans in perception and reasoning.
One core concept in multi-modal learning is how we define “multi-modality.” Some recent works suggest using task-specific definitions [Parcalabescu et al., 2021], which focus on the features and representations that are important for each task. Different tasks have different needs.This flexible view helps to guide model design more precisely. Another important topic is building taxonomies to structure the field and classify methods based on how they work and what they are used for [Baltrušaitis et al., 2018].
@@
Deep learning continuously plays a major role in multi-modal systems. It brings strong architectures that can translate one modality into another, improve representation learning, and handle multiple modalities at once [Akkus et al., 2023]. These methods have been used in many fields, such as healthcare, robotics, video understanding, and text-to-image generation. All of these show that multi-modal learning has great potential in real applications.
However, there are still many open challenges. One of them is how to design more general and powerful models that can handle unseen modalities at inference time [McKinzie et al., 2023]. Another big challenge is how to align, fuse, and reason across heterogeneous data sources in an efficient way [Liang et al., 2024c]. For large-scale use, models must also be both scalable and computationally efficient.
In recent years, researchers have increasingly focused on the robustness of multi-modal systems. Some studies propose frameworks to test or improve the stability of models under different conditions [McKinzie et al., 2023]. Robustness is essential for making models more reliable and useful in real-world tasks. At the same time, the field needs more shared definitions, taxonomies, and evaluation standards to compare methods to guide future development [Baltrušaitis et al., 2018, Parcalabescu et al., 2021].
In summary, multi-modal learning is a dynamic area with the power to change how machines interact with complex data from many sources. If we can solve problems like integration, robustness, and efficiency, for example, by optimizing models such as CLIP for low-resource situations [Akkus et al., 2023], we may unlock huge improvements in areas like autonomous driving and medical diagnosis [Liang et al., 2024c]. In the long run, progress on scalability, generalizability, and new model architectures will be the key to pushing this field forward.
@@ II Foundations of Multi-modal Processing
The foundations of multi-modal processing are rooted in the ability to integrate and handle information from different sources or modalities: text, images, audio, and video, in an effective way. Recent advancements in this field have made significant progress. Various important concepts and techniques have emerged as crucial components of multi-modal machine learning [Liang et al., 2024c]. One work proposes a taxonomy with six core technical challenges: representation, alignment, reasoning, generation, transference, and quantification [Liang et al., 2024c], demonstrating the complexity and wide range of this field.
Representation learning is a cornerstone, aiming to extract shared features across modalities to enable effective integration [Parcalabescu et al., 2021]. This process typically involves optimizing a joint loss function that balances modality-specific reconstruction with shared representation constraints.
@@
where $\mathbf{z}$ denotes the shared latent representation, $\mathbf{x}_{m}$ represents the input data from modality $m$, $\mathcal{L}_{m}$ is the modality-specific loss, $\mathcal{R}(\mathbf{z})$ enforces regularization (e.g., sparsity), and $\lambda$ balances the trade-off. This formulation ensures the model captures both modality-specific details and cross-modal commonalities. Alignment, another critical challenge, involves mapping features from different modalities into a shared space. Contrastive learning, as used in models like CLIP, is a common approach to achieve this [Akkus et al., 2023].
@@
A holistic evaluation framework is important for understanding the performance of multi-modal foundation models. Such frameworks evaluate models in three dimensions: basic abilities, information flow, and practical use in real-world scenarios [Liang et al., 2024a]. This highlights the need for better tools to measure how well multi-modal models work, and what they can really do. Meanwhile, some recent surveys have updated how we classify methods. It is no longer just about early vs. late fusion. Many works now focus on more fine-grained strategies for combining modalities [Baltrušaitis et al., 2018].
One of the major advancements is the use of task-relative definitions of multi-modality. These definitions focus on what kind of information is useful for a specific task [Parcalabescu et al., 2021]. This view supports more flexible and targeted system designs. Surveys on multi-modal co-learning also address tough problems like missing or noisy modalities. They offer new taxonomies of techniques and applications, which help guide future research [Rahate et al., 2022].
Some common themes can be found across these developments. Researchers increasingly see multi-modality as a core aspect of modern machine learning. They also recognize a set of technical and practical challenges, and propose classification systems to better structure the field. The ongoing work on evaluation frameworks, fusion strategies, and co-learning shows that this area is constantly evolving and full of new directions [Akkus et al., 2023], [Beinborn et al., 2018].
There is also growing interest in making multi-modal learning more robust. Some studies offer frameworks to analyze or improve model stability [McKinzie et al., 2023]. At the same time, the potential of multi-modal large language models (MLLMs) continues to attract attention [Wu et al., 2023]. The long-term goal is to create generalist multi-modal AI systems that can handle a wide variety of tasks and modalities, an exciting and promising direction [Munikoti et al., 2024].
However, challenges and limitations remain. Scalability is a major concern. So is the problem of missing or noisy modalities. And there is still no common standard for evaluating multi-modal models [Jabeen et al., 2023], [Shi and Niethammer, 2023]. To solve these issues, researchers are exploring methods such as correlation maximization or minimization [Shi and Niethammer, 2023], as well as improved data processing techniques for modern multi-modal architectures [Li et al., 2024c].
In conclusion, the foundations of multi-modal processing involve a rich mix of ideas, tools, and useful cases. With continued research, we can expect significant improvements in how we understand and integrate cross-modal data in higher-level models [Beinborn et al., 2018], [Wu et al., 2023]. These advances will influence many areas, like healthcare, education, entertainment, and more: underscoring the importance of continued support and funding for this field.
@@ III Deep Learning for Multi-modal Data
Deep learning has driven major progress in multi-modal research over recent years. Many novel architectures and models have been developed to integrate and process multiple types of data [Akkus et al., 2023, Suzuki and Matsuo, 2022]. A key direction in this field is multi-modal co-learning, which uses information from one modality to support another, especially when data is missing or noisy [Rahate et al., 2022]. Among these, variational autoencoders (VAEs) are often used to create multi-modal generative models that learn shared features [Suzuki and Matsuo, 2022].
@@
Architectural design for multi-modal fusion is another major research focus. For instance, EmbraceNet offers a structure that works with different types of models and avoids performance loss when some modalities are missing [Choi and Lee, 2019]. Cross-modal attention mechanisms are also important, helping align and fuse inputs from different modalities [Baltrušaitis et al., 2018]. However, scalability and computational cost remain serious concerns [McClenny et al., 2020].
Various deep learning-based methods are widely applied in this field, including neural networks, autoencoders, and generative models [Wang, 2021], [Jabeen et al., 2023]. Among these, variational autoencoders (VAEs), generative adversarial networks (GANs) and so on, are often used to create multi-modal generative models that can learn shared features and generate cross-modal outputs [Suzuki and Matsuo, 2022]. Cross-modal attention mechanisms are also important. They help align and fuse inputs from different modalities, making integration more effective [Baltrušaitis et al., 2018].
Even with these advancements, the field still faces several challenges. Scalability and computational cost remain serious concerns [McClenny et al., 2020], [Yu et al., 2020]. Handling large and diverse datasets can be expensive and complex. In addition, missing or noisy data can harm model performance, and the lack of standard architectures and evaluation benchmarks makes it hard to fairly compare different approaches. These issues show that further research is urgently needed.
Surveys of recent studies reveal several common themes: a strong focus on multi-modal fusion, the use of deep learning frameworks, concerns about robustness, and the importance of organizing the field through taxonomies and benchmarks [Summaira et al., 2021], [Thapa, 2022]. Generative models and co-learning strategies continue to attract attention. Still, researchers propose different architectures depending on task type and application needs, leading to varied designs and viewpoints.
Overall, deep learning has brought meaningful progress to multi-modal learning by enabling better integration and processing of diverse data types. However, problems like computational cost, model comparison, and noise handling remain. As the field evolves, we can expect to see more advanced models and better solutions, moving toward more effective, adaptable, and robust multi-modal systems [Menvouta et al., 2022],[Akkus et al., 2023].
@@ IV Multi-modal Fusion Methods
Fusion methods play a central and often sensitive role in multi-modal learning. They are not just technical add-ons; they are what actually make it possible to bring different modalities: text, vision, audio, or others, together in a way that works. Without good fusion, the promise of multi-modal learning becomes hard to fulfill. That’s why fusion strategies are treated as a core research focus in the community. One insight that has emerged is the importance of learning strong uni-modal features, even under supervised multi-modal training. In particular, [Du et al., 2023] introduces a kind of late-fusion learning method that aims to help models generalize better. What’s interesting here is that their approach captures the finer details unique to each modality, while also trying to limit the harm from noisy or less helpful ones. This is especially useful when the data environment is unpredictable or noisy, which is often the case in real-world settings.
@@
Early fusion combines raw features at the input level, formalized as a weighted combination of modality-specific features [Akkus et al., 2023].
@@
where $\mathbf{x}_{m}$ denotes the input features of modality $m$, $\mathbf{W}_{m}$ is a weight matrix, $\mathbf{b}$ is a bias term, and $f(\cdot)$ is a non-linear activation function (e.g., ReLU). Late fusion, conversely, integrates predictions from modality-specific models, as explored by [Du et al., 2023].
@@
where $g_{m}(\mathbf{x}_{m};\theta_{m})$ is the prediction model for modality $m$, $w_{m}$ is the modality weight, and $y_{\text{fused}}$ is the final prediction. To address modality competition, adaptive gradient modulation dynamically adjusts contributions during training [Li et al., 2023].
@@
where $\mathcal{L}_{m}$ is the loss for modality $m$, $\alpha_{m}(t)$ is a time-varying weight, and $\nabla_{\theta}\mathcal{L}$ is the fused gradient, ensuring balanced learning across modalities. Intermediate fusion methods, particularly in biomedical applications, allow gradual interaction during training [Guarrasi et al., 2024].
@@
Other researchers are taking a somewhat different path. Instead of relying purely on late-fusion, they propose models that can learn how different modalities interact and support each other, on their own. For example, the work in [Ni et al., 2023] presents a deep equilibrium model for fusion, which has shown quite impressive results on many benchmarks. The model tries to capture high-level dependencies between modalities in a very flexible way. Then, there is [Pawłowski et al., 2022], which argues that there is really no best fusion method that works for all problems. Depending on the task, the modality types, and even how much memory the system can use, the right fusion strategy may be very different.
@@
Somewhere in the middle, intermediate fusion methods offer another option. These are particularly interesting for domains like biomedical applications [Guarrasi et al., 2024], where the way signals mix can be quite subtle and context-sensitive. Intermediate fusion allows for gradual interaction between modalities during training, which often leads to better performance overall. Additionally, [Li et al., 2023] proposes a method using adaptive gradient modulation. This helps reduce what’s known as modality competition, where different data sources “compete” for influence during learning. They also suggest a new metric to measure this competition, called competition strength.
@@
Looking across these studies, there is a common theme: fusion is not just about mixing things together. It needs to be done carefully, and often in a way that adapts to both the data and the task. Some papers like [Du et al., 2023] argue for late-fusion to keep each modality’s strengths intact, while others such as [Guarrasi et al., 2024] see value in intermediate fusion, especially for structured or medical data. Adaptive methods like those in [Ni et al., 2023] and [Li et al., 2023] focus on the idea that the relationship between modalities can shift, and that the model needs to follow those shifts dynamically. Also, as [Pawłowski et al., 2022] mentions, resource availability like memory can shape what fusion strategy is even possible in practice.
@@
An issue that shows up again and again is evaluation. How do we know that fusion is working well? How do we compare models? Multiple works, including [Du et al., 2023] and [Li et al., 2023], point out the lack of widely agreed-upon metrics. Without proper tools to evaluate fusion quality or modality interaction, it is very hard to say what works and what does not.
@@
And then there is application context. Some methods are proposed for general use [Pawłowski et al., 2022], while others are more specific, like those focused on biomedical data [Guarrasi et al., 2024]. This shows the variety in this field but also points to how challenging it is to create fusion methods that are flexible yet powerful across very different tasks and domains.
@@
Even though we’ve seen a lot of progress, there are still problems that have not been solved well. Modality competition, for one, is still a bit of a black box. While [Li et al., 2023] suggests ways to measure and handle it, we do not fully understand the underlying dynamics yet. There is also the point that intermediate fusion methods, which seem promising in medical areas, have not been tested enough in other domains [Guarrasi et al., 2024]. So, we do not yet know how far they can go.
@@
EmbraceNet[Choi and Lee, 2019] 2019 Modular structure Yes Works with different model types Classification, integration
Late-fusion[Du et al., 2023] 2023 Late-fusion Not specified Captures finer details of each modality Generalization in noisy settings
Intermediate fusion[Guarrasi et al., 2024] 2024 Intermediate Yes Gradual interaction during training Biomedical applications
DEQ-Fusion[Ni et al., 2023] 2023 Deep Not specified Impressive results on benchmarks General multi-modal tasks
Adaptive Gradient[Li et al., 2023] 2023 Adaptive Not specified Reduces modality competition Balanced learning
Greedy Selection[Cheng et al., 2022] 2022 Selection-based Yes Computational efficiency Resource-constrained environments
Teacher-Student[Xue et al., 2021] 2021 Knowledge transfer Yes Works with unlabeled data Cross-modal learning
Factorized[Tsai et al., 2019] 2018 Factorized Yes Focus on shared structures Incomplete data scenarios
Asymmetric[Gao et al., 2025] 2025 Asymmetric reinforcement Yes Guides model to rely on stable signals Robust learning
Lightweight Adaptation[Reza et al., 2023] 2023 Adaptive Not specified Resource-efficient Practical applications

### v2

@@ I Introduction
Multi-modal learning, a fast-growing research field that aims to enable machines to interpret and understand the world by combining different types of data, such as images, text, and audio [2103.06304] [1705.09406]. It tries to use the strengths of each modality to build a more complete and detailed understanding of complex things [2301.04856]. This kind of learning is naturally cross-disciplinary, combining knowledge from computer vision, natural language processing, speech analysis, and other fields to make machines behave more like humans in perception and reasoning.
One core concept in multi-modal learning is how we define “multi-modality.” Some recent works suggest using task-specific definitions [2103.06304], which focus on the features and representations that are important for each task. Different tasks have different needs.This flexible view helps to guide model design more precisely. Another important topic is building taxonomies to structure the field and classify methods based on how they work and what they are used for [1705.09406].
@@
Deep learning continuously plays a major role in multi-modal systems. It brings strong architectures that can translate one modality into another, improve representation learning, and handle multiple modalities at once [2301.04856]. These methods have been used in many fields, such as healthcare, robotics, video understanding, and text-to-image generation. All of these show that multi-modal learning has great potential in real applications.
However, there are still many open challenges. One of them is how to design more general and powerful models that can handle unseen modalities at inference time [2304.04385]. Another big challenge is how to align, fuse, and reason across heterogeneous data sources in an efficient way [2209.03430]. For large-scale use, models must also be both scalable and computationally efficient.
In recent years, researchers have increasingly focused on the robustness of multi-modal systems. Some studies propose frameworks to test or improve the stability of models under different conditions [2304.04385]. Robustness is essential for making models more reliable and useful in real-world tasks. At the same time, the field needs more shared definitions, taxonomies, and evaluation standards to compare methods to guide future development [1705.09406, 2103.06304].
In summary, multi-modal learning is a dynamic area with the power to change how machines interact with complex data from many sources. If we can solve problems like integration, robustness, and efficiency, for example, by optimizing models such as CLIP for low-resource situations [2301.04856], we may unlock huge improvements in areas like autonomous driving and medical diagnosis [2209.03430]. In the long run, progress on scalability, generalizability, and new model architectures will be the key to pushing this field forward.
@@ II Foundations of Multi-modal Processing
The foundations of multi-modal processing are rooted in the ability to integrate and handle information from different sources or modalities: text, images, audio, and video, in an effective way. Recent advancements in this field have made significant progress. Various important concepts and techniques have emerged as crucial components of multi-modal machine learning [2209.03430]. One work proposes a taxonomy with six core technical challenges: representation, alignment, reasoning, generation, transference, and quantification [2209.03430], demonstrating the complexity and wide range of this field.
Representation learning is a cornerstone, aiming to extract shared features across modalities to enable effective integration [2103.06304]. This process typically involves optimizing a joint loss function that balances modality-specific reconstruction with shared representation constraints.
@@
where $\mathbf{z}$ denotes the shared latent representation, $\mathbf{x}_{m}$ represents the input data from modality $m$, $\mathcal{L}_{m}$ is the modality-specific loss, $\mathcal{R}(\mathbf{z})$ enforces regularization (e.g., sparsity), and $\lambda$ balances the trade-off. This formulation ensures the model captures both modality-specific details and cross-modal commonalities. Alignment, another critical challenge, involves mapping features from different modalities into a shared space. Contrastive learning, as used in models like CLIP, is a common approach to achieve this [2301.04856].
@@
A holistic evaluation framework is important for understanding the performance of multi-modal foundation models. Such frameworks evaluate models in three dimensions: basic abilities, information flow, and practical use in real-world scenarios [2407.03418]. This highlights the need for better tools to measure how well multi-modal models work, and what they can really do. Meanwhile, some recent surveys have updated how we classify methods. It is no longer just about early vs. late fusion. Many works now focus on more fine-grained strategies for combining modalities [1705.09406].
One of the major advancements is the use of task-relative definitions of multi-modality. These definitions focus on what kind of information is useful for a specific task [2103.06304]. This view supports more flexible and targeted system designs. Surveys on multi-modal co-learning also address tough problems like missing or noisy modalities. They offer new taxonomies of techniques and applications, which help guide future research [2107.13782].
Some common themes can be found across these developments. Researchers increasingly see multi-modality as a core aspect of modern machine learning. They also recognize a set of technical and practical challenges, and propose classification systems to better structure the field. The ongoing work on evaluation frameworks, fusion strategies, and co-learning shows that this area is constantly evolving and full of new directions [2301.04856], [1806.06371].
There is also growing interest in making multi-modal learning more robust. Some studies offer frameworks to analyze or improve model stability [2304.04385]. At the same time, the potential of multi-modal large language models (MLLMs) continues to attract attention [2311.13165]. The long-term goal is to create generalist multi-modal AI systems that can handle a wide variety of tasks and modalities, an exciting and promising direction [2406.05496].
However, challenges and limitations remain. Scalability is a major concern. So is the problem of missing or noisy modalities. And there is still no common standard for evaluating multi-modal models [2202.09195], [2305.03125]. To solve these issues, researchers are exploring methods such as correlation maximization or minimization [2305.03125], as well as improved data processing techniques for modern multi-modal architectures [2407.19180].
In conclusion, the foundations of multi-modal processing involve a rich mix of ideas, tools, and useful cases. With continued research, we can expect significant improvements in how we understand and integrate cross-modal data in higher-level models [1806.06371], [2311.13165]. These advances will influence many areas, like healthcare, education, entertainment, and more: underscoring the importance of continued support and funding for this field.
@@ III Deep Learning for Multi-modal Data
Deep learning has driven major progress in multi-modal research over recent years. Many novel architectures and models have been developed to integrate and process multiple types of data [2301.04856, 2207.02127]. A key direction in this field is multi-modal co-learning, which uses information from one modality to support another, especially when data is missing or noisy [2107.13782]. Among these, variational autoencoders (VAEs) are often used to create multi-modal generative models that learn shared features [2207.02127].
@@
Architectural design for multi-modal fusion is another major research focus. For instance, EmbraceNet offers a structure that works with different types of models and avoids performance loss when some modalities are missing [1904.09078]. Cross-modal attention mechanisms are also important, helping align and fuse inputs from different modalities [1705.09406]. However, scalability and computational cost remain serious concerns [2006.09310].
Various deep learning-based methods are widely applied in this field, including neural networks, autoencoders, and generative models [2006.08159], [2202.09195]. Among these, variational autoencoders (VAEs), generative adversarial networks (GANs) and so on, are often used to create multi-modal generative models that can learn shared features and generate cross-modal outputs [2207.02127]. Cross-modal attention mechanisms are also important. They help align and fuse inputs from different modalities, making integration more effective [1705.09406].
Even with these advancements, the field still faces several challenges. Scalability and computational cost remain serious concerns [2006.09310], [2004.12070]. Handling large and diverse datasets can be expensive and complex. In addition, missing or noisy data can harm model performance, and the lack of standard architectures and evaluation benchmarks makes it hard to fairly compare different approaches. These issues show that further research is urgently needed.
Surveys of recent studies reveal several common themes: a strong focus on multi-modal fusion, the use of deep learning frameworks, concerns about robustness, and the importance of organizing the field through taxonomies and benchmarks [2105.11087], [2211.15837]. Generative models and co-learning strategies continue to attract attention. Still, researchers propose different architectures depending on task type and application needs, leading to varied designs and viewpoints.
Overall, deep learning has brought meaningful progress to multi-modal learning by enabling better integration and processing of diverse data types. However, problems like computational cost, model comparison, and noise handling remain. As the field evolves, we can expect to see more advanced models and better solutions, moving toward more effective, adaptable, and robust multi-modal systems [2212.00101],[2301.04856].
@@ IV Multi-modal Fusion Methods
Fusion methods play a central and often sensitive role in multi-modal learning. They are not just technical add-ons; they are what actually make it possible to bring different modalities: text, vision, audio, or others, together in a way that works. Without good fusion, the promise of multi-modal learning becomes hard to fulfill. That’s why fusion strategies are treated as a core research focus in the community. One insight that has emerged is the importance of learning strong uni-modal features, even under supervised multi-modal training. In particular, [2305.01233] introduces a kind of late-fusion learning method that aims to help models generalize better. What’s interesting here is that their approach captures the finer details unique to each modality, while also trying to limit the harm from noisy or less helpful ones. This is especially useful when the data environment is unpredictable or noisy, which is often the case in real-world settings.
@@
Early fusion combines raw features at the input level, formalized as a weighted combination of modality-specific features [2301.04856].
@@
where $\mathbf{x}_{m}$ denotes the input features of modality $m$, $\mathbf{W}_{m}$ is a weight matrix, $\mathbf{b}$ is a bias term, and $f(\cdot)$ is a non-linear activation function (e.g., ReLU). Late fusion, conversely, integrates predictions from modality-specific models, as explored by [2305.01233].
@@
where $g_{m}(\mathbf{x}_{m};\theta_{m})$ is the prediction model for modality $m$, $w_{m}$ is the modality weight, and $y_{\text{fused}}$ is the final prediction. To address modality competition, adaptive gradient modulation dynamically adjusts contributions during training [2308.07686].
@@
where $\mathcal{L}_{m}$ is the loss for modality $m$, $\alpha_{m}(t)$ is a time-varying weight, and $\nabla_{\theta}\mathcal{L}$ is the fused gradient, ensuring balanced learning across modalities. Intermediate fusion methods, particularly in biomedical applications, allow gradual interaction during training [2408.02686].
@@
Other researchers are taking a somewhat different path. Instead of relying purely on late-fusion, they propose models that can learn how different modalities interact and support each other, on their own. For example, the work in [2306.16645] presents a deep equilibrium model for fusion, which has shown quite impressive results on many benchmarks. The model tries to capture high-level dependencies between modalities in a very flexible way. Then, there is [2206.06367], which argues that there is really no best fusion method that works for all problems. Depending on the task, the modality types, and even how much memory the system can use, the right fusion strategy may be very different.
@@
Somewhere in the middle, intermediate fusion methods offer another option. These are particularly interesting for domains like biomedical applications [2408.02686], where the way signals mix can be quite subtle and context-sensitive. Intermediate fusion allows for gradual interaction between modalities during training, which often leads to better performance overall. Additionally, [2308.07686] proposes a method using adaptive gradient modulation. This helps reduce what’s known as modality competition, where different data sources “compete” for influence during learning. They also suggest a new metric to measure this competition, called competition strength.
@@
Looking across these studies, there is a common theme: fusion is not just about mixing things together. It needs to be done carefully, and often in a way that adapts to both the data and the task. Some papers like [2305.01233] argue for late-fusion to keep each modality’s strengths intact, while others such as [2408.02686] see value in intermediate fusion, especially for structured or medical data. Adaptive methods like those in [2306.16645] and [2308.07686] focus on the idea that the relationship between modalities can shift, and that the model needs to follow those shifts dynamically. Also, as [2206.06367] mentions, resource availability like memory can shape what fusion strategy is even possible in practice.
@@
An issue that shows up again and again is evaluation. How do we know that fusion is working well? How do we compare models? Multiple works, including [2305.01233] and [2308.07686], point out the lack of widely agreed-upon metrics. Without proper tools to evaluate fusion quality or modality interaction, it is very hard to say what works and what does not.
@@
And then there is application context. Some methods are proposed for general use [2206.06367], while others are more specific, like those focused on biomedical data [2408.02686]. This shows the variety in this field but also points to how challenging it is to create fusion methods that are flexible yet powerful across very different tasks and domains.
@@
Even though we’ve seen a lot of progress, there are still problems that have not been solved well. Modality competition, for one, is still a bit of a black box. While [2308.07686] suggests ways to measure and handle it, we do not fully understand the underlying dynamics yet. There is also the point that intermediate fusion methods, which seem promising in medical areas, have not been tested enough in other domains [2408.02686]. So, we do not yet know how far they can go.
@@
EmbraceNet[1904.09078] 2019 Modular structure Yes Works with different model types Classification, integration
Late-fusion[2305.01233] 2023 Late-fusion Not specified Captures finer details of each modality Generalization in noisy settings
Intermediate fusion[2408.02686] 2024 Intermediate Yes Gradual interaction during training Biomedical applications
DEQ-Fusion[2306.16645] 2023 Deep Not specified Impressive results on benchmarks General multi-modal tasks
Adaptive Gradient[2308.07686] 2023 Adaptive Not specified Reduces modality competition Balanced learning
Greedy Selection[2210.12562] 2022 Selection-based Yes Computational efficiency Resource-constrained environments
Teacher-Student[2103.14431] 2021 Knowledge transfer Yes Works with unlabeled data Cross-modal learning
Factorized[1806.06176] 2018 Factorized Yes Focus on shared structures Incomplete data scenarios
Asymmetric[2501.01240] 2025 Asymmetric reinforcement Yes Guides model to rely on stable signals Robust learning
Lightweight Adaptation[2310.03986] 2023 Adaptive Not specified Resource-efficient Practical applications